# RL-based Query Rewriting with Distilled LLM for online E-Commerce Systems

### v1

@@ Abstract
RL-based Query Rewriting with Distilled LLM for online E-Commerce Systems
Duy A. Nguyen, Rishi Kesav Mohan, Van Yang, Pritom Saha Akash, Kevin Chen-Chuan Chang University of Illinois, Urbana-Champaign duyan2/rkmohan2/sy65/pakash2/kcchang@illinois.edu
@@ 1 Introduction
Previous literature. Query rewriting (QR) methods can be broadly categorized into discriminative and generative approaches. Further details about existing work are provided in Appendix A.
@@ 3.1 Offline Training Phase
The offline phase serves as a warm-start mechanism for the query rewriting model, delivers a model designed to be both (1) highly effective for the query rewriting task and (2) computationally efficient with minimal overhead.
We identify a key limitation of vanilla LLMs in the e-commerce QR task: they often generate long-tail rewrites (as shown in Table 1), which are suboptimal and difficult to process in subsequent stages of the search pipeline Peng et al. (2024); Zhang et al. (2021). To address this issue, we begin by applying supervised fine-tuning (SFT) to LLMs using a custom-built Q2Q dataset derived from the Amazon ESCI dataset Reddy et al. (2022), ensuring the models are better suited for query rewriting. Our approach trains two model variants: a Teacher model ($T$), a large-scale LLM with strong language understanding, and a Student model ($S$), a smaller, more efficient version optimized for reduced computational overhead.
Subsequently, a KD strategy is applied to transfer the Teacher model’s knowledge to the Student model, enabling $S$ to inherit the Teacher’s performance capabilities while maintaining efficiency. The result of this offline training phase is the fine-tuned Student model, aptly named MiniELM, which serves as the foundation for the subsequent online phase.
@@
SFT with Custom Q2Q dataset. We construct a custom query-to-query (Q2Q) dataset using existing queries from the Amazon ESCI dataset Reddy et al. (2022). The ESCI dataset represents data as triplets ($Q,P,R$), where $Q$ is a user query, $P$ is a product in the Amazon catalog, and $R$ is the relevance score between them. Leveraging this structure, we create a bipartite graph that maps the relevance relationships between the query set and the product set (Illustrated in Fig. 3). From this graph, we identify query pairs that are mapped as relevant to at least $k$ similar products (labeled as “Exact” or “Substitute” in the ESCI dataset). These query pairs are treated as candidate equivalents. To ensure semantic accuracy, the final set of candidate query pairs is filtered using a strong LLM (Llama 3.3, 70B version in our case), which verifies the semantic equivalence of the queries. Verified pairs are then included in the custom Q2Q dataset. This building procedure is especially beneficial, as it completely remove human manual annotations out of the loop, unlike existing works Agrawal et al. (2023); Peng et al. (2024). Building on this curated Q2Q dataset, we fine-tune both the teacher model $T$ and the student model $S$ on these query-to-query pairs. This targeted fine-tuning process ensures that both models are aligned with the task of generating accurate query rewritings within the e-commerce context. By focusing on equivalence in query rewriting, this method significantly mitigates the long-tail queries generated by vanilla LLMs.
@@
After the process, we attain the fine-tuned and distilled version of Student $S$ MiniELM, that effectively addresses the computational inefficiencies associated with deploying large-scale LLMs in real-time search systems, while maintaining great language ability and sense for E-Commerce QR task.
@@ 3.2 Online Training Phase
The online training phase extends the offline foundation by enabling MiniELM to adapt dynamically to the evolving e-commerce environment through real-time learning during deployment process (Figure 2(b)). This phase employs reinforcement learning (RL) to fine-tune the model using gradient policy optimization Schulman et al. (2017); Rafailov et al. (2024), ensuring that MiniELM remains responsive to updates in product catalogs and user behavior.
Online reward signal. The design of the metrics defined in Section 2 reflects the multifaceted goals of query rewriting. The relevance score ensures alignment between the original and rewritten queries, maintaining consistency with users’ original intents. Diversity measures the extent to which the rewritten query expands product coverage by retrieving distinct items compared to the original query. While both metrics can be calculated offline and provide a baseline reward signal, they fail to capture user interest in the retrieved products - a critical indicator of query quality. To address this, an online feedback score is derived from simulated user interactions using a judge model - named $M_{2}$. This score, combined with relevance and diversity, ensures the model balances query expansion with relevance to user preferences and broader exploratory needs. All of these metrics are quantified as follow.
@@
  • Click/Add2cart/Purchase rate score $c(P_{\tilde{Q}})/a2c(P_{\tilde{Q}})/p(P_{\tilde{Q}})$: An LLM judge model $M_{2}$ is carefully prompted to assess the quality of a rewritten query based on its associated product list $P_{\tilde{Q}}$ (detailed prompts are covered in Appendix B). The model takes as input the simulated user’s bio information, drawn from a pre-synthesized profile pool (details on the pool generation process are provided in Appendix B), along with the original query $Q$ and the product list $P_{\tilde{Q}}$. It then simulates up to $k$ interactions that the user might perform with the products. User interactions are categorized into three levels of increasing interest: clicking ($c(P_{\tilde{Q}})$), adding to cart ($a2c(P_{\tilde{Q}})$) and purchasing ($p(P_{\tilde{Q}})$). For each product list $P_{\tilde{Q}}$, $M_{2}$ is prompted to separately predict the number of interactions for each category. For instance, $c(P_{\tilde{Q}})=\frac{M2(bio,Q,P_{\tilde{Q}})}{|P_{\tilde{Q}}|}$ estimates the number of products clicked, normalized by total number of products in the list. Ideally, the interaction count should reflect the quality of $P_{\tilde{Q}}$, where higher-quality rewrites yield more positive user interactions.
@@ 4.2.2 Comparison with existing baselines
    Supervised Raffel et al. (2020): The T5 model is trained in a supervised manner with standard beam search for inference, serving as the foundational baseline for evaluating other methods.
@@
    DRQR Wang et al. (2020): An RL-based method using a reward function combining F1 score and Query Performance Predictor (QPP).
@@ 4.3 Additional Analysis
In this analysis, we delve deeper into the evolution of MiniELM’s performance and the quality of its generated queries over time during the online phase. Figure 4 illustrates the evolution of reward signals during the online training phase using DPO for both accepted and rejected rewritten queries. The queries are generated using the MiniELM model variant with a GPT2 backbone. To capture the underlying trends more effectively, the reward signals are smoothed using a mean average smoothing technique with a window size of $5$. As shown, both rewritten queries generated by MiniELM for each user query consistently improve over time, reflected in the increasing reward scores. This improvement highlights the effectiveness and consistency of our RL training process, demonstrating the model’s ability to utilize feedback from LLMs (acting as simulated human evaluators) to refine query rewritings and enhance overall performance.
@@ 5 Conclusion
This paper presents a hybrid pipeline for query rewriting in e-commerce, addressing latency, cost, and adaptability challenges. The MiniELM model balances performance and efficiency through offline knowledge distillation and online reinforcement learning. Experiments show improvements in query relevance, diversity, and user engagement. By using simulated interactions via LLMs, MiniELM adapts to evolving user preferences and catalogs, eliminating resource-intensive annotations and providing a scalable, cost-effective solution for dynamic e-commerce environments.
@@
Appendix A Related Works
A.1 Discriminative Method
A.2 Generative Method
LLMs, as demonstrated by Jagerman et al. Jagerman et al. (2023), generate semantically rich, diverse query expansions through strategies like zero-shot, few-shot, and Chain-of-Thought prompting. PRF-enhanced prompts further improve contextual understanding, but these models face challenges in fine-tuning for specific goals and impose high resource demands. Product-agent systems, such as those by Zhou et al. Zhou et al. (2024), extend LLM capabilities by integrating APIs and knowledge graphs, enabling dynamic query adaptation and addressing standalone LLM limitations.
Appendix B Prompts for Human Simulation and AI Feedback Labeling
@@
User simluation Simulate the behavior of a random e-commerce user with specific demographics and preferences influencing product choices:

### v2

# MiniELM: A Lightweight and Adaptive Query Rewriting Framework for E-Commerce Search Optimization

@@ Abstract
MiniELM: A Lightweight and Adaptive Query Rewriting Framework for E-Commerce Search Optimization
Duy A. Nguyen1∗†, Rishi Kesav Mohan1∗, Van Yang1∗, Pritom Saha Akash1∗, Kevin Chen-Chuan Chang1∗ 1Siebel School of Computing and Data Science University of Illinois, Urbana-Champaign, US VinUni-Illinois Smart Health Center, VinUniversity, Vietnam {duyan2,rkmohan2,sy65,pakash2,kcchang}@illinois.edu
@@ 1 Introduction
Previous literature. Query rewriting (QR) methods can be broadly categorized into discriminative and generative approaches. Further details about existing work are provided in Appendix 2.
@@
2 Related Works
2.1 Discriminative Method
2.2 Generative Method
LLMs, as demonstrated in Jagerman et al. (2023); Dai et al. (2024); Guo et al. (2025); Wilson et al. (2025), generate semantically rich, diverse query expansions through strategies like zero-shot, few-shot, and Chain-of-Thought prompting. PRF-enhanced prompts further improve contextual understanding, but these models face challenges in fine-tuning for specific goals and impose high resource demands. Product-agent systems, such as those by Zhou et al. Zhou et al. (2024), extend LLM capabilities by integrating APIs and knowledge graphs, enabling dynamic query adaptation and addressing standalone LLM limitations.
@@ 4.1 Offline Training Phase
The offline phase serves as a warm-start mechanism for the query rewriting (QR) model, ensuring that it is both highly effective in rewriting queries and computationally efficient with minimal overhead.
A key challenge in applying vanilla LLMs to e-commerce QR is their tendency to generate long-tail rewrites (as shown in Table 1), which are often suboptimal and difficult to process in downstream search pipeline stages Peng et al. (2024); Zhang et al. (2021). To mitigate this issue, we first apply supervised fine-tuning (SFT) using a curated Q2Q dataset derived from the Amazon ESCI dataset Reddy et al. (2022). This step adapts the model to the QR task, aligning its outputs with domain-specific requirements and improving rewrite quality. Our approach trains two model variants: a Teacher model ($T$), a large-scale LLM with strong language understanding, and a Student model ($S$), a smaller, more efficient version optimized for reduced computational overhead.
Subsequently, a KD strategy is applied to transfer the Teacher model’s knowledge to the Student model. This two-step process - fine-tuning and distillation - ensures that the Student model inherits the Teacher’s strong performance while maintaining efficiency. Fine-tuning first allows the Teacher to learn optimal QR patterns, which are then distilled into the smaller model, preventing excessive performance degradation during compression. The outcome of this offline training phase is MiniELM, a fine-tuned and distilled Student model that forms the foundation for the subsequent online phase.
@@
SFT with Custom Q2Q dataset. We construct a custom query-to-query (Q2Q) dataset using existing queries from the Amazon ESCI dataset Reddy et al. (2022). The ESCI dataset represents data as triplets ($Q,P,R$), where $Q$ is a user query, $P$ is a product in the Amazon catalog, and $R$ is the relevance score between them. Leveraging this structure, we create a bipartite graph that maps the relevance relationships between the query set and the product set (illustrated in Fig. 3). From this graph, we identify query pairs that are mapped as relevant to at least $k$ similar products (e.g. labeled as “E” or “S” in the ESCI dataset). These query pairs are treated as candidate equivalents. To ensure semantic accuracy, the final set of candidate query pairs is filtered using a strong LLM (Llama 3.3, 70B version in our case), which verifies the semantic equivalence of the queries. Final selections (e.g. “men shoes” and “shoes for men” - Figure 3) are then included in the custom Q2Q dataset. This building procedure is beneficial as it completely remove human manual annotations out of the loop, unlike existing works Agrawal et al. (2023); Peng et al. (2024). Building on this curated Q2Q dataset, we fine-tune both the teacher model $T$ and the student model $S$ on these query-to-query pairs. This targeted fine-tuning process ensures that both models are aligned with the task of generating accurate query rewritings within the e-commerce context. By focusing on equivalence in query rewriting, this method significantly mitigates the long-tail queries generated by vanilla LLMs.
@@
After the process, we attain the fine-tuned and distilled version of Student $S$ MiniELM, that addresses the computational inefficiencies associated with deploying large-scale LLMs in real-time search systems, while maintaining great language ability and sense for E-Commerce QR task.
@@ 4.2 Online Training Phase
The online training phase extends the offline foundation by enabling MiniELM to adapt dynamically to the e-commerce environment through real-time learning during deployment process (Figure 2(b)). This phase employs reinforcement learning (RL) to fine-tune the model using gradient policy optimization Schulman et al. (2017); Rafailov et al. (2024), ensuring that MiniELM remains responsive to updates in product catalogs and user behavior.
Online reward signal. To effectively guide this real-time learning, the online training phase relies on carefully designed reward signals (defined in Section 3), which capture the multifaceted objectives of query rewriting. The relevance score ensures alignment between the original and rewritten queries, maintaining consistency with users’ original intents. Diversity measures the extent to which the rewritten query expands product coverage by retrieving distinct items compared to the original query. While both metrics can be calculated offline and provide a baseline reward signal, they fail to capture user interest in the retrieved products - a critical indicator of query quality. To address this, an online feedback score is derived from simulated user interactions using a judge model - named $M_{2}$. This score, combined with relevance and diversity, ensures the model balances query expansion with relevance to user preferences and broader exploratory needs. All of these metrics are quantified as follow.
@@
  • Click/Add2cart/Purchase rate score $c(P_{\tilde{Q}})/a2c(P_{\tilde{Q}})/p(P_{\tilde{Q}})$: An LLM judge model $M_{2}$ is carefully prompted to assess the quality of a rewritten query based on its associated product list $P_{\tilde{Q}}$ (detailed prompts are covered in Appendix A). The model takes as input the simulated user’s bio information, drawn from a pre-synthesized profile pool (details on the pool generation process are provided in Appendix A), along with the original query $Q$ and the product list $P_{\tilde{Q}}$. It then simulates up to $k$ interactions that the user might perform with the products. User interactions are categorized into three levels of increasing interest: clicking ($c(P_{\tilde{Q}})$), adding to cart ($a2c(P_{\tilde{Q}})$) and purchasing ($p(P_{\tilde{Q}})$). For each product list $P_{\tilde{Q}}$, $M_{2}$ is prompted to separately predict the number of interactions for each category. For instance, $c(P_{\tilde{Q}})=\frac{M2(bio,Q,P_{\tilde{Q}})}{|P_{\tilde{Q}}|}$ estimates the number of products clicked, normalized by total number of products in the list. Ideally, the interaction count should reflect the quality of $P_{\tilde{Q}}$, where higher-quality rewrites yield more positive user interactions.
@@ 5.2.2 Comparison with existing baselines
    Supervised Raffel et al. (2020): T5 model is supervisedly trained with standard beam search for inference, serving as the foundational baseline for evaluating other methods.
@@
    DRQR Wang et al. (2020): An RL method using a reward function combining F1 score and Query Performance Predictor (QPP).
@@ 5.3 Additional Analysis
This analysis examines MiniELM’s performance evolution and query quality during the online phase. Figure 4 illustrates the evolution of reward signals during the online training phase using DPO for both accepted and rejected rewritten queries. The queries are generated using the MiniELM model variant with a GPT2 backbone. To highlight trends, rewards are smoothed using a $5$-window mean average.
As shown, both MiniELM’s rewrites consistently improve over time, reflected in rising reward scores. This improvement highlights the effectiveness and consistency of our RL training process, demonstrating the model’s ability to utilize feedback from LLMs (acting as simulated human evaluators) to refine query rewritings and enhance overall performance.
@@ 6 Conclusion
This paper introduces MiniELM, a hybrid query rewriting pipeline for e-commerce that optimizes latency, cost, and adaptability. It balances performance and efficiency through offline knowledge distillation and online reinforcement learning. Experiments show improved query relevance, diversity, and user engagement. By leveraging LLM-simulated interactions, MiniELM adapts to evolving user behavior and catalogs without costly annotations, offering a scalable, cost-effective solution for dynamic e-commerce.
@@
Ethical Considerations and Broader Impact
MiniELM introduces improvements in query rewriting for e-commerce, but its deployment should be taken with care to avoid potential ethical concerns related to bias and transparency. Since the model learns from historical data, it may reinforce existing biases, favoring popular brands or frequently searched products while underrepresenting niche sellers. Transparency is another key concern, as users and merchants have limited visibility into how and why their queries are rewritten. Without interpretability mechanisms, MiniELM’s query modifications could lead to unintended shifts in search results, affecting user trust and seller visibility.
Despite these concerns, MiniELM has the potential for significant positive impact on e-commerce search experiences if it is correctly deployed. By bridging lexical gaps and enhancing query diversity, it improves product discoverability, allowing users to find relevant items more easily, even with ambiguous or misspelled queries. This benefits both consumers and smaller sellers, as it enables lesser-known products to surface in search results. Additionally, MiniELM’s adaptive reinforcement learning mechanism ensures that query rewrites evolve with changing trends, reducing reliance on static query expansion rules. For e-commerce platforms, this leads to better search efficiency, increased user engagement, and a more scalable approach to query understanding without costly human annotations.
Acknowledgment
This material is based upon work supported by the National Science Foundation IIS 16-19302 and IIS 16-33755, Zhejiang University ZJU Research 083650, IBM-Illinois Center for Cognitive Computing Systems Research (C3SR) and IBM-Illinois Discovery Accelerator Institute (IIDAI), grants from eBay and Microsoft Azure, UIUC OVCR CCIL Planning Grant 434S34, UIUC CSBS Small Grant 434C8U, and UIUC New Frontiers Initiative. Any opinions, findings, conclusions, or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the views of the funding agencies.
The work of Duy A. Nguyen was supported in part by a PhD fellowship from the VinUni-Illinois Smart Health Center, VinUniversity, Hanoi, Vietnam.
Appendix A Prompts for Human Simulation and AI Feedback Labeling
@@
User simulation Simulate the behavior of a random e-commerce user with specific demographics and preferences influencing product choices: